### v2

# Ab initio study of the radii of oxygen isotopes

Abstract
We present an ab initio study of the charge and matter radii of oxygen isotopes from 16O to 20O using nuclear lattice effective field theory (NLEFT) with high-fidelity N3LO chiral interactions. To efficiently address the Monte Carlo sign problem encountered in nuclear radius calculations, we introduce the partial pinhole algorithm, significantly reducing statistical uncertainties and extending the reach to more neutron-rich and proton-rich isotopes. Our computed charge radii for 16O, 17O, and 18O closely match experimental data, and we predict a charge radius of $2.810(32)$ fm for 20O. The calculated matter radii show excellent agreement with values extracted from low-energy proton and electron elastic scattering data, but are inconsistent with those derived from interaction cross sections and charge-changing cross section measurements. These discrepancies highlight model-dependent ambiguities in the experimental extraction methods of matter radii and underscore the value of precise theoretical benchmarks from NLEFT calculations.
today
I Introduction
The nuclear radius, which characterizes the size of the nucleus, is one of the most fundamental properties of atomic nuclei. With the advancement of experimental techniques, there has been a rapid increase in available data on nuclear charge and matter radii for both stable and unstable isotopes Ozawa:2000gx ; Angeli:2013epw ; Campbell:2016yxe ; Yang:2022wbl . On the one hand, these data have significantly enhanced our understanding of nuclear structure, revealing novel phenomena such as nuclear halos Tanihata:1985psr ; Kanungo:2016tmz , neutron skins PREX:2021umo ; CREX:2022kgg , and nuclear shape phase transitions Marsh:2018wxs ; Cubiss:2023cyb . On the other hand, they pose considerable challenges to nuclear theory, particularly to ab initio approaches.
Despite the remarkable progress in nuclear theory, achieving accurate descriptions of nuclear charge radii remains a significant challenge Cipollone:2014hfa ; GarciaRuiz:2016ohj ; Lapoux:2016exf ; Miller:2019proton ; Koszorus:2020mgn ; Kaur:2022yoh ; Konig:2023bzp . A notable example is the oxygen isotope chain. Most ab initio calculations underestimate the charge radius of 16O by roughly 7–15% Cipollone:2014hfa ; Ekstrom:2015rta . Furthermore, while 16O and 17O exhibit similar charge radii, a pronounced increase in the charge radius is observed experimentally for 18O in comparison to 16O and 17Angeli:2013epw . To our knowledge, this intriguing feature is not captured by any existing theoretical approach. Although the chiral interaction ${\rm NNLO}_{\rm sat}$ successfully reproduces the charge radius of 16O by including it among the fitted observables Ekstrom:2015rta , it still fails to give the substantial increase seen in 18Lapoux:2016exf ; Kaur:2022yoh .
Nuclear lattice effective field theory (NLEFT) Lee:2008fa ; Lahde2019book offers a different perspective on this problem. Recently, the wave function matching method Elhatisari:2022zrb was proposed to mitigate the notorious Monte Carlo sign problem in the imaginary-time evolution of quantum many-body simulations, which introduces strong cancellations between positive and negative amplitudes. In combination with high-fidelity N3LO chiral interactions, the binding energies of both light and medium-mass nuclei up to $A=58$ are well reproduced. While charge radii are generally well described, including the pronounced increase in 18O compared to 16,17O, they are often accompanied by significantly larger statistical uncertainties. This is largely due to the commonly used pinhole algorithm, which inserts an $A$-body density operator during the imaginary-time evolution Elhatisari:2017eno . This insertion leads to unpaired nucleons which results in additional strong sign-oscillations, particularly for nuclei with larger mass numbers and/or pronounced proton–neutron asymmetry, thereby complicating the investigation of subtle isotopic trends in nuclear radii.
In this work, we propose an extension of the pinhole algorithm, referred to as the partial pinhole algorithm, that reduces the rank of the many-body density operator in order to mitigate the sign problem in nuclear radius calculations within the framework of NLEFT. This reduced operator is implemented via the recently developed rank-operator method Ma:2023ahg . We combine the partial pinhole algorithm with the wave function matching method using high-fidelity N3LO chiral interactions. As a first application, we investigate both charge and matter radii of a number of oxygen isotopes.
II Formalism
The full details of our lattice calculations are shown in the Supplemental Material SM . To illustrate the idea of the partial pinhole algorithm and how it is combined with NLEFT in calculating nuclear radii, we adopt a simplified Hamiltonian with an attractive two-body contact short range interaction ($C<0$),
$H=K+\frac{1}{2}C\,\sum_{\bm{n}}:\rho^{2}(\bm{n}):,$ (1)
where $K$ is the kinetic energy term with nucleon mass $m=938.9$ MeV, the colons indicate normal ordering, and $\rho(\bm{n})=\sum_{\sigma\tau}\rho_{\sigma\tau}(\bm{n})=\sum_{\sigma\tau}a^{\dagger}_{\sigma\tau}(\bm{n})a_{\sigma\tau}(\bm{n})$. Here, $\rho_{\sigma\tau}(\bm{n})$ is the one-body density operator at lattice site $\bm{n}$ for spin $\sigma$ and isospin $\tau$.
The ground state wave functions of $H$ can be obtained by applying imaginary time projectors to a trial wave function $\Psi_{0}$, $|\Psi\rangle=\lim_{\tau\to\infty}e^{-H\tau/2}|\Psi_{0}\rangle$. In Monte Carlo simulations, $\tau$ is divided into $L_{t}$ slices with temporal spacing $a_{t}$ such that $\tau=L_{t}a_{t}$. For each time slice, the two-body interaction is defined as nucleons propagated in a fluctuating background auxiliary field using a Hubbard-Stratonovich transformation,
$\exp\left(-\frac{a_{t}C}{2}\rho^{2}\right)=\sqrt{\frac{1}{2\pi}}\int ds~\exp\left(-\frac{s^{2}}{2}+\sqrt{-a_{t}C}s\rho\right),$ (2)
where $s$ is the auxiliary field at a lattice site. Therefore, the wave function $|\Psi\rangle$ can be written as an auxiliary field path-integral expression,
$|\Psi\rangle=\int\mathcal{D}{s_{1}}\cdots\mathcal{D}{s_{L_{t}/2}}M(s_{Lt/2})\cdots M(s_{1})|\Psi_{0}\rangle~,$ (3)
where the initial wave function $|\Psi_{0}\rangle$ is a $A$-nucleon Slater determinant, such as alpha clusters states or shell-model wave functions, and the normal-ordered transfer matrix $M$ is defined as,
$M(s_{n_{t}})=:\exp\left[-a_{t}K+\sqrt{-a_{t}C}\sum_{\bm{n}}s_{n_{t}}(\bm{n})\rho(\bm{n})\right]:.$ (4)
Therefore, $|\Psi\rangle$ is expressed as a linear combination of $A$-nucleon Slater determinants, each of which is associated with a specific configuration of auxiliary fields $\vec{s}$.
With the wave function $|\Psi\rangle$ in Eq. (3), the root-mean-square (RMS) point-proton radius of a nucleus with proton number $Z$, neutron number $N$, and mass number $A$ can be evaluated by
$\langle r_{p}^{2}\rangle=\frac{1}{Z}\sum_{\bm{n},\sigma,\tau=p}\langle\rho_{\sigma\tau}(\bm{n})(\bm{n}-\hat{\bm{r}}_{\rm c.m.})^{2}\rangle,$ (5)
where $\hat{\bm{r}}_{\rm c.m.}=\sum_{\bm{n},\sigma,\tau}\bm{n}\rho_{\sigma\tau}(\bm{n})/A$ is the center-of-mass operator. The expression can be rewritten in terms of a set of two-body correlation functions $G_{\alpha\beta}$ as
$\langle r_{p}^{2}\rangle=\frac{N}{ZA^{2}}G_{pp}+\frac{N-Z}{ZA^{2}}G_{pn}-\frac{1}{A^{2}}G_{nn}~,$ (6)
with
$G_{\alpha\beta}=\sum_{\begin{subarray}{c}\bm{n}_{1},\sigma_{1},\tau_{1}=\alpha\\ \bm{n}_{2},\sigma_{2},\tau_{2}=\beta\end{subarray}}\langle:\rho_{\sigma_{1}\tau_{1}}(\bm{n}_{1})\rho_{\sigma_{2}\tau_{2}}(\bm{n}_{2}):\rangle(\bm{n}_{1}-\bm{n}_{2})^{2}.$ (7)
Similarly, the matter radius is given by
$\langle r_{m}^{2}\rangle$ $=\frac{1}{A}\sum_{\bm{n},\sigma,\tau}\langle\rho_{\sigma\tau}(\bm{n})(\bm{n}-\hat{\bm{r}}_{\rm c.m.})^{2}\rangle$
$=\frac{1}{2A^{2}}G_{pp}+\frac{1}{A^{2}}G_{pn}+\frac{1}{2A^{2}}G_{nn}.$ (8)
Thus, nuclear radii can be expressed in terms of just three two-body correlation functions. It is important to note that since the matter radius corresponds to the second radial moment of the intrinsic nuclear density distribution (or equivalently, the squared nuclear wave function), its numerical value is sensitive to both the theoretical framework used and the specific methods adopted when extracting it from experimental observables.
Similar to the previous studies on computing the RMS charge radii Ekstrom:2015rta ; Elhatisari:2022zrb ; Sun:2025yfo , we use the standard relation Friar:1975Advances
$r_{\rm ch}^{2}=\langle r_{p}^{2}\rangle+R_{p}^{2}+\frac{N}{Z}R_{n}^{2}+\frac{3}{4m_{p}^{2}},$ (9)
where $R_{p}^{2}=0.7056~{\rm fm}^{2}$ Pohl:2010zza ; Lin:2021xrc , $R_{n}^{2}=-0.105~{\rm fm}^{2}$ Filin:2020tcs , and $m_{p}=938.27$ MeV. Relativistic spin-orbit corrections and two-nucleon current contributions are not included in this work.
A direct calculation of the correlation functions $G_{\alpha\beta}$ is computationally expensive, especially when the corrections from high-fidelity chiral interactions are included perturbatively. To evaluate them more efficiently, we introduce the partial pinhole algorithm. We define a normal-ordered $M$-body density operator,
$\rho_{M}(\vec{c}\,)=:\rho_{\sigma_{1}\tau_{1}}(\bm{n}_{1})\cdots\rho_{\sigma_{M}\tau_{M}}(\bm{n}_{M}):,$ (10)
using the notation $\vec{c}=(c_{1},\cdots,c_{M})$. Here, $c_{i}=(\bm{n}_{i},\sigma_{i},\tau_{i})$ specifies the quantum numbers of the $i$-th density operator, with $\bm{n}_{i}$ denoting the lattice coordinate, and $\sigma_{i}$ and $\tau_{i}$ representing spin and isospin. The summation over $\vec{c}$ satisfies
$\sum_{\vec{c}}\rho_{M}(\vec{c}\,)$ $=\hat{N}(\hat{N}-1)\cdots(\hat{N}-M+1)$
$=A(A-1)\cdots(A-M+1),$ (11)
where $\hat{N}$ is the particle-number operator. The equality of the second line stems from the fact that we are working in $A$-nucleon subspace. Therefore, this summation is proportional to the identity operator.
The correlation functions $G_{\alpha\beta}$ can then be evaluated as
$G_{\alpha\beta}$ $=\frac{A(A-1)}{M(M-1)}$
$\quad\times\frac{\displaystyle\sum_{\vec{c}}\left[\langle\Psi|\rho_{M}(\vec{c}\,)|\Psi\rangle\sum_{i<j}\delta_{\tau_{i}\alpha}\delta_{\tau_{j}\beta}(\bm{n}_{i}-\bm{n}_{j})^{2}\right]}{\displaystyle\sum_{\vec{c}}\langle\Psi|\rho_{M}(\vec{c}\,)|\Psi\rangle}.$ (12)
The details about computing $(\bm{n}_{i}-\bm{n}_{j})^{2}$ are presented in the Supplemental Material SM . Unlike the standard pinhole algorithm Elhatisari:2017eno , the value of $M$ is much smaller than the total number of nucleons $A$. As a result, $\rho_{M}$ is no longer a projection operator, and acting with $\rho_{M}$ on a single $A$-nucleon Slater determinant yields a linear combination of multiple $A$-nucleon Slater determinants.
Since $\rho_{\sigma\tau}(\bm{n})$ is a rank-one operator, that is, $\rho_{\sigma\tau}^{2}(\bm{n}):=0$, the rank-one operator method Ma:2023ahg allows us to express $\rho_{M}(\vec{c}\,)$ as
$\rho_{M}(\vec{c}\,)=\frac{1}{\varepsilon^{M}}:\exp[\varepsilon\rho_{\sigma_{1}\tau_{1}}(\bm{n}_{1})+\cdots+\varepsilon\rho_{\sigma_{M}\tau_{M}}(\bm{n}_{M})]:$ (13)
with $\varepsilon\to\infty$. It consists of a string of one-body operators which can be applied directly to each single-particle wave function in the Slater determinant.
The summations over $\vec{c}$ in Eq.(II) and the path integral over $\vec{s}$ in Eq.(3) are evaluated via Monte Carlo importance sampling. We generate an ensemble of $\vec{c},\vec{s}$} configurations according to the relative probability weight,
$P(\vec{c},\vec{s}\,)=|\langle\Psi_{0}|$ $M(s_{L_{t}})\cdots M(s_{L_{t}/2+1})\rho_{M}(\vec{c}\,)$
$\times M(s_{L_{t}/2})\cdots M(s_{1})|\Psi_{0}\rangle|.$ (14)
The strings of transfer matrices are generated by updating the auxiliary fields $\vec{s}$ using the shuttle algorithm Lu:2018bat . At the middle time slice, the auxiliary field updates are paused, and the pinhole configuration $\vec{c}$ is updated by performing standard Metropolis accept/reject steps. Specifically, to update $\vec{c}$, we randomly select an index $i$ in $\rho_{M}$ and either move it to a neighboring site,
$c_{i}=\{\bm{n}_{i},\sigma_{i},\tau_{i}\}\rightarrow c_{i}^{\prime}=\{\bm{n}_{i}^{\prime},\sigma_{i},\tau_{i}\},$ (15)
or reassign its spin and isospin,
$c_{i}=\{\bm{n}_{i},\sigma_{i},\tau_{i}\}\rightarrow c_{i}^{\prime}=\{\bm{n}_{i},\sigma_{i}^{\prime},\tau_{i}^{\prime}\}.$ (16)
The new configuration $\vec{c}\,^{\prime}$ is accepted if
${P(\vec{c}\,^{\prime},\vec{s}\,)}/{P(\vec{c},\vec{s}\,)}>r,$ (17)
where the random number $r$ is uniformly distributed between 0 and 1.
In this work, we combine the partial pinhole algorithm with the wave function matching method. The original chiral Hamiltonian is unitarily transformed into a new high-fidelity Hamiltonian $H$, whose wave function matches that of a computationally simple Hamiltonian, $H_{S}$, at a given radius. This ensures rapid convergence of the perturbative expansion in $H-H_{S}$. The details can be found in Ref. Elhatisari:2022zrb . The formalism for applying the partial pinhole algorithm in perturbative calculations is provided in SM .
III Results and discussion
In the following calculations, we adopt the N3LO chiral interactions from Ref. Elhatisari:2022zrb . We also use a minimal sets of three-body force terms, where, in addition to nuclear binding energies, the charge radius of 4He is used to determine the low-energy constants (LECs) Elhatisari:pre . Although the number of involved free LECs is reduced from eight to six, the overall description of nuclear binding energies is similar to Ref. Elhatisari:2022zrb , as detailed in SM . Our simulations are performed with a spatial lattice spacing of $a=1.32~{\rm fm}$, corresponding to a momentum cutoff $\Lambda=\pi/a\approx 471~{\rm MeV}$. Additionally, we use a temporal lattice spacing of $a_{t}=0.001~{\rm MeV}^{-1}$ and a cubic box of length $L=13.2~{\rm fm}$. Since the radii of the oxygen isotopes under investigation are all below 3 fm, this box size is clearly sufficient to suppress finite-volume effects.
In the partial pinhole algorithm, although one could in principle take $M=2$ for $\rho_{M}$ in Eq. (10) as the correlation functions $G_{\alpha\beta}$ in Eq. (7) are two-body observables, such a choice suffers from low computational efficiency. To improve sampling efficiency, we aim for each configuration $\vec{c}$ to contribute simultaneously to all three types of correlations: $G_{pp}$, $G_{pn}$, and $G_{nn}$. This requirement implies that the isospin indices in $\rho_{M}$ must include at least two protons and two neutrons. We therefore adopt $M=4$ and fix half of the indices to be protons and the other half to be neutrons in our calculations. We have further verified that the final results are largely insensitive to the specific choice of $M$, having tested $M=4$, 6, 8, and 10 SM .
In SM , we benchmark the charge radii of 16O and 17O using both the standard pinhole and the partial pinhole algorithms with the simple Hamiltonian $H_{S}$. Although we find that both methods yield consistent charge radii, the partial pinhole algorithm maintains a significantly larger average phase factor and substantially reduces statistical uncertainties, particularly in the case of 17O. This improvement becomes even more substantial for heavier nuclei such as 40Ca, and is further amplified when applying the full N3LO chiral interactions.
[…]
Table 1: The charge and matter radii (units are fm) calculated by the partial pinhole algorithm using the N3LO interaction, where the theoretical uncertainties are estimated using ten sets of three-nucleon forces. The available experiment data are also given for comparison. Exp. $(e,e)$ Angeli:2013epw : charge radii deduced from electron scattering data. Exp. $(e,e)$, $(p,p)$ Lapoux:2016exf : matter radii deduced from electron scattering and proton scattering data. Exp. $\sigma_{I}$ Ozawa:2000gx : matter radii deduced from interaction cross sections. Exp. $\sigma_{cc}$ Kaur:2022yoh : matter radii deduced from charge changing cross sections.
16O 17O 18O 20O
$r_{\rm ch}$ [NLEFT] $2.704(17)$ $2.709(15)$ $2.768(17)$ $2.810(32)$
Exp. $(e,e)$ $2.699(5)$ $2.693(8)$ $2.776(2)$
$r_{m}$ [NLEFT] $2.576(17)$ $2.651(14)$ $2.744(19)$ $2.863(33)$
Exp. $(e,e)$, $(p,p)$ $2.60(8)$ $2.67(10)$ $2.77(10)$ $2.90(10)$
Exp. $\sigma_{I}$ $2.54(2)$ $2.59(5)$ $2.61(8)$ $2.69(3)$
Exp. $\sigma_{cc}$ $2.57(2)$ $2.64(8)$ $2.71(3)$
[…]
The extrapolated charge radii are summarized in Table 1, where results using nine additional sets of three-nucleon forces are included to estimate theoretical uncertainties Elhatisari:pre . Note that the central values and the corresponding errors are shifted compared to the extrapolated ones in Fig. 1. The results show that 16O and 17O have very similar charge radii, while 18O exhibits a noticeable increase, consistent with experimental data Angeli:2013epw . We note that, although 16O is commonly described as a spherical doubly magic nucleus within shell-model-based approaches, NLEFT simulations predict a distinct tetrahedral arrangement of alpha clusters Epelbaum:2013paa . Similar alpha-cluster structures in 8Be Shen:2024qzi , 12Shen:2022bak , and 20Ne Harris:2025zsh have also been predicted by NLEFT in their ground states. Supporting this picture, previous NLEFT studies Shen:2022bak demonstrated that the ground state of 12C robustly converges to an equilateral triangular arrangement of alpha clusters independent of the choice of the initial state. By analogy, assuming that 16O similarly forms a stable tetrahedral cluster configuration as indicated by NLEFT calculations, this intrinsic clustering could explain the success of NLEFT in reproducing oxygen isotope charge radii.
We also compute the matter radii of oxygen isotopes using the partial pinhole algorithm. The results for 16O, 17O, 18O, and 20O are shown in Fig. 2. Similar to the charge radii shown in Fig. 1, we perform a double-exponential fit to extrapolate the data to $\tau\to\infty$; see SM for details. In contrast to the charge radii, the matter radii increase monotonically with mass number. The extrapolated matter radii are listed in Table 1, along with values obtained using nine additional sets of three-nucleon forces Elhatisari:pre .
In Table 1, we also compare our results with experimental matter radii extracted from electron and proton scattering [$(e,e)$ and (p,p)(p,p)Lapoux:2016exf , interaction cross sections ($\sigma_{I}$Ozawa:2000gx , and charge-exchange cross sections ($\sigma_{cc}$Kaur:2022yoh . Unlike charge radii, the extraction of matter radii often involves some degree of model dependence, leading to discrepancies among different experimental methods. For example, matter radii derived from low-energy proton elastic scattering tend to be systematically larger than those extracted from $\sigma_{I}$ and $\sigma_{cc}$ measurements. The most notable case is 20O, where differences up to 0.2 fm are observed among the three different measurements. As discussed in Ref. Lapoux:2016exf , the $(e,e)$ and $(p,p)$ data provide a more reliable descriptions, since the correlations in the target are usually not included in the other methods. From Table 1, we can find our calculated matter radii agree well with the values extracted from $(e,e)$ and $(p,p)$ data, and thus support the statement given in Ref. Lapoux:2016exf .
IV Summary and perspectives
We have performed an ab initio calculation of the charge and matter radii of oxygen isotopes from 16O to 20O using nuclear lattice effective field theory (NLEFT) with the high-fidelity N3LO chiral interactions. To mitigate the Monte Carlo sign problem in nuclear radius calculations, we introduce the partial pinhole algorithm, which reduces the rank of the many-body density operator in the standard pinhole approach and significantly lowers the statistical uncertainties. This method enables access to much larger projection times and extends the applicability to more neutron-rich and proton-rich nuclei. Our results accurately reproduce the experimental charge radii of 16O, 17O, and 18O, demonstrating the capability of NLEFT in describing nuclear structure properties. The charge radius of 20O is predicted to be $2.810(32)$ fm. The calculated matter radii show excellent agreement with values extracted from low-energy electron and proton scattering, while deviating from those inferred from interaction and charge-exchange cross sections. These results underscore the need for further experimental clarification, particularly for neutron-rich isotopes, with improved precision and reduced model dependence.
As a general and versatile method, the partial pinhole algorithm holds broad potential future applications. It can be integrated with other Monte Carlo techniques, such as the perturbative quantum Monte-Carlo method Lu:2021tab , and has already been successfully applied to the calculation of additional observables, including electromagnetic transitions Shen:2024qzi . Its ability to suppress sign oscillations while maintaining high accuracy suggests wide applicability not only across nuclear systems, but also to broader classes of quantum many-body problems.
Acknowledgements.
We are grateful for discussions with Fabian Hildenbrand, Timo Lähde, Dean Lee, Bing-Nan Lu, Yuan-Zhuo Ma, Xiang-Xiang Sun, and Shuang Zhang. This work was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101018170), and by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). The work of SE is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK project no. 123F464). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). Furthermore, the authors gratefully acknowledge the computing time provided on the high-performance computer HoreKa by the National High-Performance Computing Center at KIT (NHR@KIT). This center is jointly supported by the Federal Ministry of Education and Research and the Ministry of Science, Research and the Arts of Baden-Württemberg, as part of the National High-Performance Computing (NHR) joint funding program (https://www.nhr-verein.de/en/our-partners). HoreKa is partly funded by the German Research Foundation (DFG).